OpenStudy (anonymous):

150 workers were engaged to finish a job in a certain number of days. 4 workers dropped out on second day, 4 more workers dropped out on third day and so on. It took 8 more days to finish the work. Find the number of days in which the work was completed.

OpenStudy (hba):

Let x be the number of days in which 150 workers finish the work.

OpenStudy (hba):

Are you familiar with A.P ?

OpenStudy (hba):

@Aditi_Singh

OpenStudy (anonymous):

\[\frac{ n(2*150+(n-1)*-4) }{ 2 } =150(n-8)\]

OpenStudy (anonymous):

Find n..)

OpenStudy (anonymous):

Yep, ofcourse i know AP!

OpenStudy (mayankdevnani):

Let the work be completed in n days when 4 workers are dropped on every day except on the first day. Since 4 workers are dropped on every day except on the first day, total number of man days who worked for n days is the sum of n terms of an A.P. with first term as 150 and common difference as –4. The work would have finished in (n − 8) days with 150 workers when the number of workers are not drooped. Total number of man days who would have worked at all n days = 150(n – 8) ∴n (152 – 2n) = 150(n – 8) ⇒152n – 2n 2 = 150n – 1200 ⇒ 2n 2 –2n – 1200 = 0 ⇒ n 2 – n – 600 = 0 ⇒ (n – 25)(n + 24) = 0 ⇒ n = 25 [n > 0] Thus, the number days in which the work was to be completed is (25 – 8) = 17 days

OpenStudy (mayankdevnani):

ok @Aditi_Singh

OpenStudy (mayankdevnani):

got it @Aditi_Singh

OpenStudy (anonymous):

no @mayankdevnani :(

OpenStudy (anonymous):

I don't get that n-8 wala part!

OpenStudy (mayankdevnani):

answer is wrong

OpenStudy (anonymous):

No, it's absolutely right!

OpenStudy (mayankdevnani):

ok @Aditi_Singh

OpenStudy (anonymous):

@Yahoo! , How do you get tha n-8 part ? @hba , Help me friend!

OpenStudy (anonymous):

the question said it took 8 more Days....with 150 , 146 ...etc Workers on 1 st 2 2nd .....Days..... so if 150 Workers are working it would take n-8 days

OpenStudy (hba):

so 17+8=25